OpenStudy (anonymous):

(2x+3)/(x+2)^4 by what method do i integrate this??

OpenStudy (anonymous):

break 2x + 3= 2(x+2)/(x+2)^4 - 1/(x+2)^4 = 2/(x+2)^3 -1/(x+2)^4 now you can integrate easily